I am having trouble solving these problems. Could you kindly help me.
1. y+1/2=-1/3(x+1/2)
2. What was the MSRP for a Hummer that sold for $ 107,272 after an 8% discount?

Thank you kindly for helping me.

Answer by stanbon(48558)

1. y+1/2=-1/3(x+1/2)
y = (-1/3)x - 1/3 - 1/2
y = (-1/3)x -[1/3 + 1/2]
y = (-1/3)x - (5/6
---------------------------
2. What was the MSRP for a Hummer that sold for $ 107,272 after an 8% discount?
Let the MSRP be "x"
x - 0.08x = 107,272
0.92x = 107,272
x = $116,600
================
Cheers,
Stan H.